OpenStudy (anonymous):

find the distance between (0,0) and (-2,-2).round that to the nearest tenth

OpenStudy (anonymous):

magintude? sqrt(2^2+2^2) sqrt(8) ... 2.8.. i guess.. weirdly worded question

OpenStudy (anonymous):

\[Distance = \sqrt{\Delta y^2 + \Delta x^2}\] \[Distance = \sqrt{(0-(-2))^2+(0-(-2))^2}\] \[Distance = \sqrt{(2)^2+(2)^2}\] \[Distance = \sqrt{4+4}\] \[Distance = \sqrt{8}=\sqrt{2*4}=2\sqrt{2}=2.8\]